OpenStudy (shamil98):

\[\int_{}^{} \sin^3 2x~~dx\] so uh how do i do this?

OpenStudy (shamil98):

\[\int_{}^{} (sin^2 2x) * (sin 2x) dx \] \[\int_{}^{} 1-cos^2 2x * \sin 2x ~dx \] i got this far..

OpenStudy (sidsiddhartha):

use \[\sin3x=3sinx-4\sin^3x\\sin^3x=\frac{ 3sinx-\sin3x }{ 4 }\] well known identity

OpenStudy (sidsiddhartha):

so \[\sin^3(2x)=\frac{ 3\sin(2x)-\sin(6x) }{ 4 }\]

OpenStudy (shamil98):

oh.

OpenStudy (sidsiddhartha):

\[\int\limits_{}\sin^3(2x)=\frac{ 3 }{ 4 }\int\limits_{}\sin(2x)dx-\frac{ 1 }{ 4 }\int\limits_{}\sin(6x)dx\] easy now :)

OpenStudy (shamil98):

yeah got it from here, tnx

OpenStudy (sidsiddhartha):

welcome :)